OpenStudy (anonymous):

The hypotenuse of an isosceles right triangle is 7 feet longer than either of its legs. Find the exact length of each side.

OpenStudy (anonymous):

a^2 + b^2 = c^2 a = b c = a+7 = b+7 a^2 + a^2 = (a+7)^2

OpenStudy (p0sitr0n):

x = hypotenuse x-7 = side 1 and 2 Pythagorean theorem 2(x-7)^2 -x^2 = 0

OpenStudy (p0sitr0n):

i think this is pretty easy to be asked for on the site.

OpenStudy (anonymous):

hypo - leg = 7 √2x - x = 7 => x = 16.9 => hypo = 16.9√2 = 23.9